OpenStudy (anonymous):

Find the sum of the geometric series –5, -10, -20, -40, . . . for 7 terms.

OpenStudy (anonymous):

I suppose the answer is -635 can any1 reassure me

OpenStudy (kropot72):

The formula needed is: \[S _{n}=\frac{a(r ^{n}-1)}{r-1}\] Where a is the first term and r is the common ratio. So plugging the values into the formula: \[S _{n}=\frac{-5(2^{7}-1)}{2-1}=you\ can\ calculate\]

OpenStudy (anonymous):

is it - 635??

OpenStudy (anonymous):

I calculated it

OpenStudy (anonymous):

@kropot72

OpenStudy (kropot72):

Correct!

OpenStudy (anonymous):

thanks

OpenStudy (kropot72):

You're welcome :)